Systems Staff Meeting Notes 5/11/98

Systems Staff Meeting Notes
May 11, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Bobbie Parker, Suzy Shaw, Leota Whitfield

Announcements
  1. Phones - voice mail feature to be available soon. Barbara is negotiating with PPD. Useable feature synopsis: park (if line is NOT busy), camp (if line IS busy), messages -feature 1(to leave), feature 65#(to delete), feature 871-to transfer ringing to all phones, page 1-1 will get you the intercom for internal speakers.
  2. Pagers - Systems has received 5 pagers to be used by: Bobbie, Lee, Winston, Bill & pseudo Destry. The numbers are listed under the Systems folder on NT2. The numbers are to be used by Systems personnel only (Bill will give his number out for weekend/PM emergencies). These were purchased at a flat rate - personal use is fine. Leaves numeric messages only - leave the phone number & your extension.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Lee has heard rumors that there are incoming printers & Cows on the loading dock this morning.
  2. One laptop for MSL was received.
  3. The modem was delivered to MSL - it is now in the conference room.
  4. Two low-end Toshiba laptops are on order for Special Collections.
Networking
  1. NT3 is back on line.
  2. Office ’97 is on NT2 for local installations. Liaisons are to keep track of the number of installations they make.
  3. Outlook ’98 will be the next step.
  4. We are removing any trace of AppleTalk cables. Will also be removing all Farallon hubs.
  5. West & Education are scheduled for the switching hub installation on May 19. The closet in West has been cleaned out by Lee & Bobbie & the hub will be housed there.
  6. Lee moved the Circulation SNA’s & printer on the Control Unit. Once the rest are done, we’ll be able to pull the Control Unit.
  7. Need to do wire abatement in West.
  8. As a result of the switch, IP’s and gateways in Education will need to be changed. West will only need to change gateways.
  9. If the IP block is stable, we’ll need to get Circulation access defined. Also need to find a scanner stand.
  10. If the switch experiment works properly, we will continue installing more - East the next week, then MSL. AFA & Music will come later - JM will be left as is.
Software Projects Status
  1. Mellon is finished - Winston has David’s final version.
  2. Sharon is still working on Personnel migration.
  3. OPS tracking is on hold.
  4. ILL will convert to their new invoicing system on June 5. Lending side only. Reporting will be done with CrystalReports.
  5. Brian is working on timer software for public work stations.
DP Coordinator
  1. Stephanie & Becky are the new students starting today - Jason will be returning.
  2. Scott will work today, he is still unsure of his schedule.
  3. Student schedules are posted on SELMA.
Calendar
  1. Debra out May 15.
  2. Brian out May 21-22.
  3. Memorial Day observed May 25.
  4. Bobbie out June 8 - 19.
  5. Debra out half day June 16, all day June 17.
  6. Bill & Winston out June 15-24.
  7. Independence Day observed July 3.
  8. Debra out July 7 - 10.
  9. Bobbie out July 13 - 17
  10. Brian out July 16 & 17.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu